OpenStudy (anonymous):

A town's public works department has ordered sand to spread on roadways during the upcoming winter snows. The delivered sand forms a right cone 27 feet high and a base diameter of 48 feet. How many cubic yards of sand has the department received?

OpenStudy (anonymous):

Well, remember that the volume of a cone is given by the formula: \[V = \frac{ 1 }{ 3 } \pi r^2 h\] And that radius (r) = D/2. So you have radius and height. Plug in the equation, what do you get?

OpenStudy (anonymous):

I got 16286.01632 feet